OpenStudy (anonymous):

please just someone tell me how to do this !!! i need to get this over and done with plzzzz!!! :( 1.Using complete sentences, define and compare radian measure to degree measure. In doing so, be sure to answer each of the following questions: When would degree measure be appropriate? When would radians be a better choice? What are the pros and cons of each? 2. Adam and Beth are visiting an amusement park and have decided to ride the carousel. Adam picked a horse on the outside edge and Beth chose a dragon on the inside, closer to the center.

OpenStudy (anonymous):

Part 1: Do Adam and Beth travel the same distance during the ride? Choose a distance that each seat (horse and dragon) sits from the center and use the radius to determine how far each would travel during one rotation. Part 2: Choose an angle of rotation. Using complete sentences, compare the distance Adam and Beth will travel during this angle measurement. Part 3: Using complete sentences, describe which position you would prefer and why.

OpenStudy (anonymous):

3.robbie the robot is on a weather satellite orbiting earth about 3600 km above the surface.the earths radius is about 6400 km .he has had a malfunction in his output device and the satellite is traveling without communication . his last report was only in terms of trigonometric values and was only partially received it said , \[\sin \theta <0\] and then he was lost again Part 1: Create a set of coordinates that would be reasonable for Robbie’s position in space and satisfy his last, partial report. Using complete sentences, describe Robbie’s location and your reasoning. Part 2: What are the values of the \[\sin \theta \cos \theta and \tan \theta \] using your coordinate point?